OpenStudy (anonymous):

Use a transformation to solve the equation. x – 0.7 = 1.9 A. x = –2.6 B. x = –1.2 C. x = 1.2 D. x = 2.6 s

OpenStudy (ryanwhite499):

Let's solve your equation step-by-step. x−0.7=1.9 Step 1: Add 0.7 to both sides. x−0.7+0.7=1.9+0.7 x=2.6 Answer: x=2.6 @fallengoddess

OpenStudy (anonymous):

so its d

OpenStudy (ryanwhite499):

Yes :)

OpenStudy (ryanwhite499):

Close this question